OpenStudy (anonymous):

A spherical object has a diameter of 18 cm. It has a spherical inner core with a diameter of 9 cm. What is the volume of the other layer? Use 3.14 to approximate pi and express your answer in hundredths. _____cm3

OpenStudy (campbell_st):

so the radius of the object is 9 cm the radius of the core is 4.5 cm then using the formula for a sphere \[V = \frac{4}{3} \times \pi \times r^3\] the volume of the outer layer = Volume of large sphere - volume of inner core hope it helps

OpenStudy (anonymous):

Okay thank you :) It helps, I had forgot the formula lol

OpenStudy (anonymous):

I got 2670.57 is that correct?

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

yes that's what I got

OpenStudy (anonymous):

Ok thank you but it says express answer in hundredths